# Investigating Energy Efficiency and Performance Trade-offs in LLM Inference Across Tasks and DVFS Settings

### v2

@@ III-A1 Energy Consumption
Mistral-7B consistently exhibits the highest energy consumption across all tasks, ranging from approximately 2,800 to 3,100 joules per 1,200 queries (Fig 1). In contrast, LLaMA-3.2-1B operates with remarkable energy efficiency, consuming only about 650–730 joules for the same query load, a 77% reduction compared to Mistral-7B. Falcon-7B also demonstrates better energy efficiency than Mistral, averaging around 2,650–2,900 joules. GPT-Neo-2.7B is notably inefficient; despite having fewer parameters than Falcon and Mistral, it consumes up to 6,968 joules on BoolQ which is more than double the energy used by models with higher performance. This highlights that architectural inefficiencies of such larger unoptimized KV caches and slower attention mechanisms that can dominate over parameter count in dictating energy cost.

### v3

@@ III-A1 Energy Consumption
Figure 1 reports per-query energy costs across models and tasks. Mistral-7B consumes between 2,800 and 3,100 J for 1,200 queries, placing it among the most energy-intensive models, as expected for its scale. In contrast, LLaMA-3.2-1B is the most efficient, requiring only 650 to 730 J for the same workload, which is about a 77% reduction relative to Mistral-7B. Falcon-7B shows slightly better efficiency than Mistral, averaging 2,650 to 2,900 J, though it remains far more costly than LLaMA-3.2-1B.
The most striking case is GPT-Neo-2.7B. Despite having fewer parameters than Falcon-7B and Mistral-7B, it consumes up to 6,968 J on BoolQ, which is more than twice the energy of the larger models. This anomaly suggests that architectural inefficiencies such as unoptimized key–value cache handling, slower attention mechanisms, and memory bandwidth bottlenecks can outweigh parameter count as the dominant factor in energy consumption.